# 20-22 December 2024, Cox’s Bazar, Bangladesh

### v1

@@ Abstract
The increasing demand for underwater exploration and rescue operations enforces the development of advanced wireless or semi-wireless underwater vessels equipped with manipulator arms. This paper presents the implementation of a semi-wireless underwater vehicle, ”TritonZ” equipped with a manipulator arm, tailored for effective underwater exploration and rescue operations. The vehicle’s compact design enables deployment in different submarine surroundings, addressing the need for wireless systems capable of navigating challenging underwater terrains. The manipulator arm can interact with the environment, allowing the robot to perform sophisticated tasks during exploration and rescue missions in emergency situations. TritonZ is equipped with various sensors such as Pi-Camera, Humidity, and Temperature sensors to send real-time environmental data. Our underwater vehicle controlled using a customized remote controller can navigate efficiently in the water where Pi-Camera enables live streaming of the surroundings. Motion control and video capture are performed simultaneously using this camera. The manipulator arm is designed to perform various tasks, similar to grasping, manipulating, and collecting underwater objects. Experimental results shows the efficacy of the proposed remotely operated vehicle in performing a variety of underwater exploration and rescue tasks. Additionally, the results show that TritonZ can maintain an average of 13.5cm/s with a minimal delay of 2-3 seconds. Furthermore, the vehicle can sustain waves underwater by maintaining its position as well as average velocity. The full project details and source code can be accessed at this link:
https://github.com/kawser-ahmed-byte/TritonZ
@@
 [^1] This is the author’s version of the work accepted at IEEE ICCIT 2024. © 2024 IEEE. Personal use permitted. Final version at https://doi.org/10.1109/ICCIT64611.2024.11022145
@@ I Introduction
Seven-tenths of the Earth’s surface is covered by oceans, yet much of this underwater world remains unexplored [1]. This lack of exploration is largely due to the challenges and risks involved in deep-sea diving and exploration [1]. Underwater robots and vehicles have emerged as vital tools for addressing these challenges, enabling technology-driven exploration. The increasing use of Remotely Operated Vehicles (ROVs) in fields such as underwater exploration, environmental monitoring, and rescue missions has driven active research into efficient, lightweight designs with multi-dimensional applications [2, 3, 4].
Recent advancements have focused on improving ROV manipulation capabilities for tasks like water sampling [5], maritime security, environmental inspection [6], and archaeological data collection [7]. Advances in underwater imaging have further enabled applications like automatic fish recognition [8] and IoT-based water quality monitoring [9].
Efforts in this field include AUVs capable of withstanding extreme conditions [10] and cost-effective wireless robots for environmental monitoring [11]. Integration of underwater manipulator arms has improved control and interaction in subsea environments [12, 13].
However, commercial underwater systems often face challenges, including low accuracy, limited automation, and restricted control capabilities due to technical and environmental constraints [14, 15]. These gaps highlight the need for compact, efficient, and accessible ROV designs.
The paper is structured as follows: Section II reviews related works, Section III focuses on the mathematical modeling, Section IV outlines the proposed ROV methodology, Section V analyzes the investigation results and discussions, and Section VI concludes the study while offering suggestions for future work.
II Related Works
Underwater robots, such as ROVs and AUVs, have transformed subsea exploration. ROVs have been used since the 1960s, while AUVs and USVs now offer autonomous access to hazardous environments, though their adoption faces technical and commercial barriers [16]. Efforts to enhance underwater safety include a rescue vehicle for drowning victims [17] and a low-cost AUV localization algorithm using acoustic signals, validated through simulations and sea trials [18].
Ali et al. [19] developed a twin-controller system combining PID and MRAC for faster convergence and reduced control error in underwater robots. Bykanova et al. [20] created a compact ROV with a rotating wing, ensuring efficient maneuverability during wreck surveys. Omerdic et al. [21] introduced a marine operation platform with 3D displays, voice navigation, and auto-tuning controllers to optimize underwater tasks.
In manipulation research, Perez and Vasquez [22] developed a 6-DOF mathematical model to improve stability and navigation, outperforming traditional systems. Other advances include precise sensor positioning for high-frequency seafloor data [23], SONAR integration for low-light conditions [24], and autonomous manipulators capable of target capturing [25].
Low-cost ROVs have also emerged, featuring battery protection, live camera feeds, and safety mechanisms for underwater exploration [26]. Another ROV focused on biofouling removal, reducing human intervention and enhancing maintenance efficiency [27]. The ”ROVer,” a lightweight hexagonal ROV for marine education, demonstrated enhanced payload capacity and effective video recording [28].
Despite advancements, commercial ROVs still struggle with low precision, limited automation, and high costs. This research presents TritonZ, a cost-effective, semi-wireless ROV with a robotic arm, 4-DOF navigation, and real-time video streaming for efficient underwater exploration and rescue missions.
III Mathematical Modeling of the Proposed ROV
@@
Mathematical modeling is fundamental for understanding the dynamics of ROVs. We wanted to maintain stability in roll and pitch direction due to the separation between their centers of gravity and buoyancy. With minimal roll and pitch motion under typical conditions, the ROV dynamics are best described by a coupled 3-degree-of-freedom (DOF) model in the XY-plane and a separate 1-DOF model in depth along the Z-axis [29]. In TritonZ’s design, four vertically aligned thrusters are used for up-down motion whereas two horizontal thrusters are for left-right-forward-backward movement. This configuration aligns well with the theoretical framework of 4 DOF models. This section focuses on the 3-DOF XY-plane model, assuming independent depth control, equivalent to a 4-DOF ROV model. The general equation of finite-dimensional models is used to represent the 4 DOFs encompassing surge, sway, and heading, which can be expressed in the following form [30],
@@ III-E Transformation Matrix
This matrix converts the ROV’s body-fixed velocities to the inertial frame, crucial for accurate navigation and control. It shows how pitch changes affect velocity, allowing us to design a control system that adapts to these variations, ensuring steady movement and precise path tracking.
III-F Rotational Matrix
@@
III-G Net Forces in the Body Frame
@@
The expression balances the ROV’s weight and buoyancy for stable vertical movement. By optimizing buoyant component placement, we achieve neutral buoyancy, reducing thruster reliance for better depth control. A comprehensive derivation of all the expressions of motion can be found in [31].
IV Proposed ROV Architecture
@@
IV-A 3D Design of TritonZ
After thorough mathematical modelling, we designed the 3D model which lays the foundation for physical implementation. The dimensions, components, and corresponding materials were aligned with our mathematical modeling and chosen carefully for underwater suitability. PVC pipe was selected for structural components due to its mechanical properties, influencing the ROV’s maximum operational depth, calculated by $P=\rho\cdot g\cdot h$. The 3D design model shown in Fig. 1 was developed using AutoDesk’s Fusion 360 [32].
@@
The ROV, as shown in Fig. 1, has dimensions of 166mm height, 705mm width, and 640mm length, resembling a quadcopter drone. It features four vertically-oriented and two horizontal thrusters, custom-designed and 3D-printed using PLA+ material. The body, made of PVC pipe, supports a 474mm manipulator arm with an aluminum gripper. Components are assembled with 4-inch PVC pipes, T-joints, and elbows.
Weight distribution is crucial for underwater performance, with the body, motors, and propellers being the main contributors. The battery and buoyancy systems also play vital roles. Designed for lightweight, easy deployment, the ROV includes a Pi camera for real-time streaming, a gripper arm for manipulation, and a custom two-blade propeller for navigation. Control is semi-wireless, using ESP modules and WiFi.
@@
IV-B Internal architecture of TritonZ
@@
IV-C System Implementation
The ROV’s control system uses wireless communication between two ESP8266 modules (operator and ROV) via the ESP-NOW protocol, enabling two-way control and sensor data transmission. The Arduino Mega handles motor control and sensor processing, while Electronic Speed Controllers (ESCs) manage motor speed and direction. The servo-controlled aluminum gripper allows precise object handling, and custom 3D-printed PLA two-blade propellers ensure efficient propulsion. The system integrates a Flask server with OpenCV and Python for live video streaming and object detection, with the Raspberry PiCam providing high-resolution underwater footage. A user-friendly HTML and CSS interface facilitates control and streaming.
@@ IV-C System Implementation
V Results & Discussion
Comprehensive experimental trials to determine the efficiency of our ROV were performed. Despite its compact size (60 cm x 70 cm x 25 cm) and lightweight (11 kg), the ROV demonstrated impressive mobility and adaptability in varied underwater environments as shown in Fig. 3(c). The integrated manipulator arm exhibited precise and agile functionality. The Pi cam’s real-time video streaming and object detection significantly enhanced situational awareness and mission accuracy. The ROV is equipped with a DHT22 sensor for monitoring water temperature and humidity, offering ±0.5°C and ±2% RH accuracy, with 98.15% reliability [33]. Additionally, the ROV features a voltage sensor with ±0.1V accuracy for real-time battery status [34]. Both sensors performed reliably during trials, ensuring accurate measurements for underwater tasks, with future versions focusing on precision improvement and additional sensor integration. Notably, the prototype surpassed runtime expectations, operating for 45 minutes using two 12.5V, 2200mAh batteries, showcasing enhanced operational capacity. Furthermore, the ROV’s affordability, with a prototype cost of approximately 12,000 BDT (110 USD), shows its potential for widespread application.
The following discussion elaborates on the results related to the three main components of our ROV:
@@ V-A Analysis of Motion
Furthermore, we also plotted the displace and velocity vs. time in still-water testing to understand how the ROV behaves with time and the time it takes to reach the average velocity. Fig. 4 (trial-1) shows the velocity-displacement profiles over a displacement of 400 cm. We observe from Fig. 4 that the ROV achieves maximum velocity within 6s. After achieving maximum velocity under still-water condition, the ROV maintains this speed till the finish line. We also tested our ROV under wave-induced conditions in the water. The results of the wave-induced condition is shown in table II (trial-4). Due to the disturbance in the water, the ROV took 39s which is higher compared to previous trials and achieved less average velocity, 12.25cm/s. Fig. 5 shows the motion in wave-induced condition. Here, the ROV also achieves maximum velocity within 8s as we can see from the figure. Wave is induced in the environment at 30s and after that we see the velocity drops and displacement along the surge direction decreases. However, our ROV has been designed keeping the disturbance of the surrounding environment. Our trials demonstrated the ROV’s capabilities in both normal and wave-induced water conditions, with consistent results observed across three tests in each scenario (Fig. 4, 5). However, broader testing in strong currents, varying water temperatures, and deeper underwater conditions is needed. Future work will focus on these challenges to provide a more comprehensive evaluation of TritonZ’s reliability and versatility. The ROV can maintain its position due to the symmetric veritical thrusters and buoyancy system even under induced wave. This can be observed from (Fig. 5) where the ROV increases its velocity within 5s and reaches around 12 cm/s to complete the overall distance.
@@
Previous research on underwater remotely operated vehicles (ROVs) has largely focused on basic features like navigation and image capture, often neglecting advanced capabilities such as object manipulation and real-time streaming. Our ROV addresses these limitations with a robotic arm for precise recovery and live-streaming functionality. Table III showcases how TritonZ outperforms existing models in transparency, control range, weight, and cost. Built with cost-effective PVC and a 1000kv BLDC motor, TritonZ is lightweight, stable, and optimized for underwater operations, minimizing speed loss and ensuring steady buoyancy. These features make it ideal for education, marine research, environmental monitoring, and rescue missions.
@@
VI Conclusion & Future Work
Our proposed system focuses on building a lightweight, efficient underwater vehicle with a cost-effective, semi-wireless, remotely controlled underwater ROV. The prototype demonstrates diverse capabilities, covering various functionalities, and rescue operations, including a Pi Cam and a manipulator arm for object interaction. The ROV can maintain reasonable speed and sustain waves and other disturbances in the water while navigating in 4 dimensions. Additionally, it can send live data to the user end such as live video footage underwater, humidity, and temperature of the water. In the future, we aim to enhance the accessibility of the ROV’s arm by incorporating more degrees of freedom, addressing the current limitation of its 4-DOF design. This improvement will allow for better handling of complex manipulation tasks, increasing the ROV’s versatility and effectiveness. Additionally, we want to enhance the performance and refine the structure of our ROV so that it can perform heavier tasks in real time.

### v2

# TritonZ: A Remotely Operated Underwater Rover with Manipulator Arm for Exploration and Rescue Operations

@@ Abstract
The increasing demand for underwater exploration and rescue operations enforces the development of advanced wireless or semi-wireless underwater vessels equipped with manipulator arms. This paper presents the implementation of a semi-wireless underwater vehicle, ”TritonZ” equipped with a manipulator arm, tailored for effective underwater exploration and rescue operations. The vehicle’s compact design enables deployment in different submarine surroundings, addressing the need for wireless systems capable of navigating challenging underwater terrains. The manipulator arm can interact with the environment, allowing the robot to perform sophisticated tasks during exploration and rescue missions in emergency situations. TritonZ is equipped with various sensors such as Pi-Camera, Humidity, and Temperature sensors to send real-time environmental data. Our underwater vehicle controlled using a customized remote controller can navigate efficiently in the water where Pi-Camera enables live streaming of the surroundings. The motion control and video capture are performed simultaneously, utilizing this camera. The manipulator arm is designed to perform various tasks, similar to grasping, manipulating, and collecting underwater objects. Experimental results shows the efficacy of the proposed remotely operated vehicle in performing a variety of underwater exploration and rescue tasks. Additionally, the results show that TritonZ can maintain an average of 13.5cm/s with a minimal delay of 2-3 seconds. Furthermore, the vehicle can sustain waves underwater by maintaining its position as well as average velocity.
@@
footnotetext: DISCLAIMER: This is the preprint version of the manuscript, as submitted before peer review. The accepted version is published in: 2024 27th International Conference on Computer and Information Technology (ICCIT), DOI: https://doi.org/10.1109/ICCIT64611.2024.11022145. © 2025 IEEE. Personal use permitted. For other uses, permission must be obtained from IEEE.
@@ I Introduction
Seven-tenths of the earth’s surface is covered by oceans, yet most of this underwater world remains unexplored [1]. This lack of exploration is largely due to the challenges and risks involved in deep-sea diving and exploration [1]. Underwater robots and vehicles are now designed for technology-driven exploration. With the increasing use of underwater remotely operated vehicles (ROVs) in various fields such as underwater exploration and analysis, commercial transportation, and rescue missions, active research is focused on designing efficient, lightweight ROVs with multi-dimensional applications [2, 3, 4, 5].
Recent works in underwater robotics include ROVs having autonomous manipulation capabilities i.e. interacting with the underwater environment, navigation, control, etc[6]. Underwater manipulation is challenging, especially in tasks involving physical interactions. Previous research includes ROVs performing water sampling and video surveillance [7], maritime security and environmental inspection [8], data collections from submerged archaeological sites to create a 3D virtual model [9]. Additionally, it has been used to precisely position sensors and equipment on the seafloor to collect high-frequency data [10]. Our research focuses on contributing to this growing field by designing and developing underwater electric manipulators with robotic arms for physical interactions with the environment [11]. Additionally, we align with existing initiatives exploring the use of enhanced underwater images which have various applications such as automatic fish recognition [12], water quality using IoT [13], etc.
In Malaysia, a study suggested a specialized design for an autonomous underwater vehicle (AUV) that can endure underwater pressures and temperatures for navigation [14]. In another study [15], the authors developed an affordable fully wireless aquatic robot. This robot was designed for tasks such as naval survey, environmental monitoring, and ash checks. Recent works have focused on Wireless Underwater ROVs incorporating underwater manipulator arms [16, 17]. Ali et al. [18] proposed a twin-controller method combining PID and MRAC to control an underwater robot’s behavior. The characteristics of the ROV were validated through experiments, showing fast convergence, reduced error, and disturbance mitigation. Bykanova et al. [19] designed a compact Remotely operated underwater vehicle (ROUV) with a variable restoring moment using a rotating wing, aimed at surveying underwater wrecks. Their proposed system maintained acceptable speed ranges for both normal (20.75 to 0.75 m/s) and longitudinal motions (21 to 1 m/s). Furthermore, Ali et al. [20] focused on position stabilization of a fully actuated ROUV with a gripper using an 8-degree-of-freedom control method. They demonstrated reduced inaccuracy and deflection, achieving the desired positioning in simulations. Omerdic et al. [21] developed a platform with advanced 3D displays, voice navigation, and auto-tuning controllers to enhance marine operations. In another study, Perez and Vasquez [22] replicated a Visor3 system, developing a 6-DOF mathematical model for underwater vehicle navigation and control. This proposed system demonstrated effective vehicle motion control, showing improved stability over conventional methods.
However, existing Commercial underwater manipulator systems exhibit limited control capabilities and automation, with low accuracy, repeatability, and control loop frequency due to environmental challenges, technical limitations, etc. [23],[24, 25]. So, there are a lot of scopes for designing more compact and efficient wireless/semi-wireless ROVs with physical interaction capabilities. In this work, we have developed a lightweight, semi-wireless ROV named ”TritonZ”, designed for efficient underwater operation, featuring surveillance and manipulation capabilities [26]. Equipped with a robotic arm that can be controlled remotely, our ROV allows for effective interaction with the environment while streaming real-time video footage with a clear view of their surroundings. It can navigate in 4 degrees of freedom (DOF) using a wireless remote, ensuring precise maneuverability for various underwater tasks. The performance of our ROV such as displacement-velocity along the surge directions, stability, controls, etc. was discussed. Additionally, its easy-to-deploy design and cost-effective approach make it accessible to a wide range of users, making it an ideal solution for underwater exploration and research.
II Mathematical Modeling of the Proposed ROV
@@
Mathematical modeling is fundamental for understanding the dynamics of ROVs. We wanted to maintain stability in roll and pitch direction due to the separation between their centers of gravity and buoyancy. With minimal roll and pitch motion under typical conditions, the ROV dynamics are best described by a coupled 3-degree-of-freedom (DOF) model in the XY-plane and a separate 1-DOF model in depth along the Z-axis [27]. In TritonZ’s design, four vertically aligned thrusters are used for up-down motion whereas two horizontal thrusters are for left right-forward-backward movement. This configuration aligns well with the theoretical framework of 4 DOF models. This section focuses on the 3-DOF XY-plane model, assuming independent depth control, equivalent to a 4-DOF ROV model. The general equation of finite-dimensional models is used to represent the 4 DOFs encompassing surge, sway, and heading, which can be expressed in the following form [28],
@@ II-E Transformation Matrix
This matrix is essential for converting the ROV’s body-fixed velocities into the inertial frame for accurate navigation and control. Our case showed how changes in the ROV’s pitch would affect its overall velocity. Understanding this allowed us to design a control system that adjusts for these changes, keeping the ROV on a steady course and accurately following its planned path, even when its orientation changes.
II-F Rotational Matrix
@@
II-G Net Forces in the Body Frame
@@
The above expression was used to balance the ROV’s weight and buoyancy for controlled vertical movement. By using the rotational matrix to analyze force distribution, we optimized the placement of buoyant materials and components to achieve neutral buoyancy. This balance helps the ROV maintain its depth efficiency, without relying heavily on thrusters. A comprehensive derivation of all the expressions of motion could be found here [29].
III Proposed ROV Architecture
@@
III-A 3D Design of TritonZ
After thorough mathematical modelling, we designed the 3D model which lays the foundation for physical implementation. The dimensions, components, and corresponding materials were aligned with our mathematical modeling and chosen carefully for underwater suitability [30, 31]. PVC pipe was selected for structural components due to its mechanical properties, influencing the ROV’s maximum operational depth, calculated by $P=\rho\cdot g\cdot h$. The 3D design model shown in Fig. 1 was developed using AutoDesk’s Fusion 360 [32].
@@
As shown in Fig. 1, the ROV has dimensions of 166mm in height, 705mm in width, and 640mm in length, closely resembling a quadcopter drone. It is equipped with four vertically oriented thrusters and two additional thrusters for horizontal movement, custom-designed by our team using PLA+ material and 3D-printed. The remainder of the ROV’s body is constructed from PVC pipe. The ROV also features a 474mm manipulator arm with a aluminum gripper. This design enables precise movement and rotation. The structure was assembled using 4-inch PVC pipes, along with T joints and elbows, to complete the ROV’s framework. The weight distribution of each ROV component is crucial to its underwater performance. As shown in table I, the main contributors to the overall weight are the body, motors, and propellers. Understanding these weights is essential for assessing buoyancy, mobility, and payload capacity. Additional elements, such as the battery and buoyancy systems, also play a significant role. Our ROV is designed to be lightweight and drone-like for easy deployment, featuring a Pi camera for real-time video streaming and a gripper arm for manipulation tasks. Navigation is enhanced by a custom two-blade propeller, with semi-wireless control achieved via ESP modules and WiFi communication.
@@
III-B Internal architecture of TritonZ
@@
III-C System Implementation
The ROV’s control system utilizes a custom wireless communication setup between an ESP8266 module on the operator’s side and another on the ROV, both programmed using Arduino IDE. The ESP-NOW protocol facilitates two-way communication for sending control commands and receiving sensor data. The ESP 8266 microcontrollers enable wireless communication between the user controller and the ROV, allowing for remote control and data transmission. The Arduino Mega handles motor control, processes sensor data, and coordinates with the ESP8266. Electronic Speed Controllers (ESCs) manage the speed and direction of the brushless motors, ensuring precise and safe operation. The aluminum gripper, controlled by a servo motor, allows for accurate object handling underwater. Custom-made two-blade, 3D-printed in PLA, deliver efficient underwater propulsion, enhancing the ROV’s movement. The system’s software includes remote control features for the Raspberry Pi, enhancing exploration capabilities. A custom-designed Flask server, utilizing OpenCV and Python, manages live video streaming and object detection. The Raspberry PiCam Module paired with the flask server provides high-resolution video for real-time underwater monitoring and object detection. A user-friendly interface developed using HTML and CSS provides real-time streaming and control functionalities.
@@ III-C System Implementation
IV Results & Discussion
Comprehensive experimental trials to determine the efficiency of our ROV were performed. Despite its compact size (60 cm x 70 cm x 25 cm) and lightweight (11 kg), the ROV demonstrated impressive mobility and adaptability in varied underwater environments as shown in Fig. 3(c). The integrated manipulator arm exhibited precise and agile functionality. The Pi cam’s real-time video streaming and object detection significantly enhanced situational awareness and mission accuracy. Notably, the prototype surpassed runtime expectations, operating for 45 minutes using two 12.5V, 2200mAh batteries, showcasing enhanced operational capacity. Furthermore, the ROV’s affordability, with a prototype cost of approximately 12,000 BDT (110 USD), shows its potential for widespread application. The following discussion elaborates on the results related to the three main components of our ROV:
@@ IV-A Analysis of Motion
Furthermore, we also plotted the displace and velocity vs. time in still-water testing to understand how the ROV behaves with time and the time it takes to reach the average velocity. Fig. 4 (trial-1) shows the velocity-displacement profiles over a displacement of 400 cm. We observe from Fig. 4 that the ROV achieves maximum velocity within 6s. After achieving maximum velocity under still-water condition, the ROV maintains this speed till the finish line. We also tested our ROV under wave-induced conditions in the water. The results of the wave-induced condition is shown in table II (trial-4). Due to the disturbance in the water, the ROV took 39s which is higher compared to previous trials and achieved less average velocity, 12.25cm/s. Fig. 5 shows the motion in wave-induced condition. Here, the ROV also achieves maximum velocity within 8s as we can see from the figure. Wave is induced in the environment at 30s and after that we see the velocity drops and displacement along the surge direction decreases. However, our ROV has been designed keeping the disturbance of the surrounding environment. The ROV can maintain its position due to the symmetric veritical thrusters and buoyancy system even under induced wave. This can be observed from 5 where the ROV increases its velocity within 5s and reaches around 12 cm/s to complete the overall distance.
@@
Previous research on underwater remotely operated vehicles (ROVs) has mainly focused on basic features like navigation and capturing images. However, these studies often lack advanced capabilities such as object manipulation and live-streaming. Our ROV addresses these gaps by including a robotic arm for precise object recovery and real-time video streaming. Table III highlights how our ROV differs from existing models by offering full transparency in aspects like control range, weight, and cost. Using cost-effective materials such as PVC for the frame and a 1000kv BLDC motor, we maintain affordability while enhancing functionality. These choices make our ROV suitable not only for educational use but also for marine research, environmental monitoring, search and rescue missions, and inspecting underwater structures.
@@
V Conclusion
Our proposed system focuses on building a lightweight, efficient underwater vehicle with a cost-effective, semi-wireless, remotely controlled underwater ROV. The prototype demonstrates diverse capabilities, covering various functionalities, and rescue operations, including a Pi Cam and a manipulator arm for object interaction. The ROV can maintain reasonable speed and sustain waves and other disturbances in the water while navigating in 4 dimensions. Additionally, it can send live data to the user end such as live video footage underwater, humidity, and temperature of the water. In this way, our ROV can send different sensor data and have a wide range of applicability for real-life tasks. In the future, we aim to enhance the accessibility of the ROV’s arm by providing at least 3 degrees of freedom for our gripper. Additionally, we want to enhance the performance and refine the structure of our ROV so that it can perform heavier tasks in real time.